### v1

@@ Abstract
With the rapid growth of artificial intelligence (AI) in healthcare, there has been a significant increase in the generation and storage of sensitive medical data. This abundance of data, in turn, has propelled the advancement of medical AI technologies. However, concerns about unauthorized data exploitation, such as training commercial AI models, often deter researchers from making their invaluable datasets publicly available. In response to the need to protect this hard-to-collect data while still encouraging medical institutions to share it, one promising solution is to introduce imperceptible noise into the data. This method aims to safeguard the data against unauthorized training by inducing degradation in model generalization. Although existing methods have shown commendable data protection capabilities in general domains, they tend to fall short when applied to biomedical data, mainly due to their failure to account for the sparse nature of medical images. To address this problem, we propose the Sparsity-Aware Local Masking (SALM) method, a novel approach that selectively perturbs significant pixel regions rather than the entire image as previous strategies have done. This simple-yet-effective approach significantly reduces the perturbation search space by concentrating on local regions, thereby improving both the efficiency and effectiveness of data protection for biomedical datasets characterized by sparse features. Besides, we have demonstrated that SALM maintains the essential characteristics of the data, ensuring its clinical utility remains uncompromised. Our extensive experiments across various datasets and model architectures demonstrate that SALM effectively prevents unauthorized training of deep-learning models and outperforms previous state-of-the-art data protection methods.
Keywords:
Unlearnable Example Unauthorized Training Data Protection Adversarial Machine Learning Classification Segmentation.
@@ 1 Introduction
The expansion of artificial intelligence (AI) in healthcare has significantly increased the production and storage of sensitive medical data [50, 52]. This data plays a crucial role in advancing research in related fields [68, 32, 52]. The more high-quality, open-source datasets that are available, the more contributions can be made by talented researchers to the development of the field [51, 29, 27, 48, 7, 52]. However, many dataset creators are reluctant to open-source their work due to concerns over unauthorized use, such as training commercial AI models.
@@
Conventional image classification datasets, which are relatively easy to obtain and label by laypersons, as shown by ImageNet’s use of online crowdsourcing [12, 70]. But in the medical field, annotating is a much more complex process requiring specialized knowledge, usually undertaken by experts like radiologists [33, 58, 30]. In addition to identifying the type of disease, detailed annotations regarding the severity of the disease and its relation to surrounding tissues are also necessary. This means that constructing a medical dataset is a labor-intensive task. Unauthorized use could lead to infringement of the creator’s rights, leading to a reluctance to publicly release further data. Moreover, from a patient’s perspective, concerns about their information being exploited for commercial purposes might decrease their willingness to authorize their data for research [34, 61]. Therefore, the construction and release of a medical dataset are not only time-consuming and labor-intensive but also fraught with significant ethical and privacy challenges. The reduction in quality open source datasets, resulting from both of the scenarios mentioned, in turn, slows down the development of medical AI [2, 14, 19].
@@
To defend unauthorized use and encourage sharing data, [25] proposed the technique of contaminating data with imperceptible noise. Models trained on this noise-contaminated data exhibit poor performance for normal utilization. Specifically, these methods create “unlearnable” examples from clean data by adding imperceptible, error-minimizing noise, demonstrating significant protective capabilities. The design of this error-minimizing noise is based on the intuitive idea that an example with a higher training loss could contain more information to learn. Consequently, this noise protects the data by minimizing the corresponding loss, effectively reducing the informativeness of the data.
@@
Nevertheless, applying this method directly to medical images may not be optimal, as it overlooks the unique properties inherent in medical images [42]. The most important property that is overlooked is sparsity [65, 10, 26, 47, 11, 13]. For instance, in cellular microscopy images, even after cropping, there remains a substantial amount of blank background. Similarly, techniques such as CT or tomographic scanning inherently produce sparse data [8, 11, 13]. Previous methods often struggled to pinpoint specific feature regions in medical data, inadvertently emphasizing sparse areas when generating noise. This not only resulted in a significant waste of computational resources but also rendered the noise performance suboptimal, potentially making it ineffective in the workflow of medical data and significantly delaying the open-source release of the data.
@@
    To address these issues, we propose Sparsity-Aware Local Masking (SALM), which is specifically designed for the medical domain by limiting the perturbation to features improved protection effectiveness and robustness.
@@
    Experiments on multiple medical datasets from MedMNIST across different modalities, scales, and tasks demonstrate that our SALM achieves SoTA performance and is consistently effective in various medical scenarios. Additionally, we validate the flexible applicability of SALM in real-world situations through case studies.
@@ 2 Related Works
2.1 Unlearnable Example
In this paper, we seek to protect data from unauthorized exploitation via a data poisoning approach. Data Poisoning is a technique used to compromise the performance of machine learning models on clean data by deliberately altering training samples. This form of attack has proven effective against both Deep Neural Networks (DNNs) and traditional machine learning methods such as SVM [6]. [45] has highlighted the susceptibility of DNNs to data poisoning, although these attacks typically result in only a modest reduction in DNN performance. However, [62] found there is a clear distinction between poisoned samples and normal samples, which comes at the cost of reduced data usability. Backdoor Attacks represent a specialized form of data poisoning. Traditional backdoor attacks [9, 43] involve the introduction of falsely labeled training samples embedded with covert triggers into the dataset. A relatively new approach within this realm is the creation of Unlearnable Examples [25]. These are considered a more subtle form of backdoor attack, free of labels and triggers. Unlearnable Examples show promising results in protecting data from unauthorized exploitation in various domains and applications [41, 40, 39, 60, 69, 36, 21, 71, 54, 20], such as natural language processing [28], graph learning [39], and contrastive learning [53]. However, the flexibility of protecting medical data has not yet been fully explored. Recently, [42] took a first step to evaluate the performance of conventional image protection methods on medical images. Nevertheless, the methods in [42] are straightforward adaptations from the previous approach and are suboptimal due to a lack of consideration of the intrinsic characteristics of medical data. To better address this, in this study, we leverage the inherent features of medical data to design more effective protection.
2.2 Medical Data Protection
As information technology advances, digital technologies are increasingly integrated into medicine, affecting both clinical treatment and scientific research [55, 51, 29, 27, 48, 7, 52, 68, 32]. However, these developments also present significant risks and challenges regarding patient privacy, medical information breaches occur frequently around the world [3, 31, 49, 46]. Inadequate data protection can lead to substantial harm through the leakage of personal information, such as the disclosure of sensitive health conditions like HIV, potentially leading to social isolation and psychological disorders [18]. Furthermore, incidents of data breaches may also reduce patients’ trust in medical research institutions, making them reluctant to share their data [34, 61]. Earlier studies focused on robust physical protection measures, like using encrypted storage devices [23, 1, 59], establishing firewalls [5, 37], and secure communication transmission modes [17, 35], to safeguard data. [38] introduced a clustering method based on $K$-anonymity algorithm as the building block of privacy-preserving for medical devices’ data. However, as collaboration among various research institutions intensifies and open-source data sharing on the internet becomes more crucial, these methods become less applicable. [4, 66] attempted to use generated data to reduce granularity and thus protect privacy, but this method involves a significant trade-off between information loss and protection efficacy. However, our method remains robust under various conditions during the construction process of real-world datasets and effectively balances high data usability with consistent protection.
@@ 3.1 Problem Statement
Suppose the clean training dataset consists of $n$ clean examples, that is, $\mathcal{D}_{c}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{n}$ with $\boldsymbol{x}\in\mathcal{X}\subset\mathbb{R}^{d}$ are the inputs and $y\in\mathcal{Y}=\{1,\cdots,K\}$ are the labels and $K$ is the total number of classes. We denote its unlearnable version by $\mathcal{D}_{u}=\left\{\left(\boldsymbol{x}_{i}^{\prime},y_{i}\right)\right\}_{i=1}^{n}$, where $\boldsymbol{x}^{\prime}=\boldsymbol{x}+\boldsymbol{\delta}$ is the unlearnable version of training example $\boldsymbol{x}\in\mathcal{D}_{c}$ ,and $\boldsymbol{\delta}\in\Delta\subset\mathbb{R}^{d}$ is the “invisible” noise that makes $\boldsymbol{x}$ unlearnable. The noise $\boldsymbol{\delta}$ is bounded by $\|\delta\|_{p}\leq\epsilon$ with $\|\cdot\|_{p}$($L_{p}$ norm), and $\epsilon$ is set to be small such that it does not affect the normal utility of the example.
@@
3.2 Error-Minimizing Noise
In an ideal scenario, the generation of noise involves a class-matching process to protect each category within $\mathcal{D}_{c}$. For simplicity, we define the noise generation process on $\mathcal{D}_{c}$ here. Given a clean example $\boldsymbol{x}$, by generating the least disruptive noise $\boldsymbol{\Delta}$ for the training input $\boldsymbol{x}$ by solving the following bi-level optimization problem:
@@ 3.2 Error-Minimizing Noise
Input: Training data set $\mathcal{T}$, Training steps $M$, PGD parameters $\alpha_{u}$ and $\rho_{u}$, transformation distribution $T$, the sampling number $J$ for gradient approximation
1 Initialize source model parameter $\theta$.
2 Randomly initialize $\delta^{u}$
3 // Following Eq. (3)
4 for $i$ in $1,\cdots,M$ do
5       Sample minibatch $(x,y)\sim\mathcal{T}$, sample transformation $t_{j}\sim T$
6       Calculate $g_{k}\leftarrow\frac{1}{J}\sum_{j=1}^{J}\frac{\partial}{\partial\delta^{u}}\ell(f^{\prime}_{\theta}(t_{j}(x+\delta^{u})),y)$
7       Determine $k$-th percentile value $g^{0}(k)$ in $g_{k}$
8       for each element $g_{ij}$ in $g_{k}$ do
9             if $g_{ij}\geq g^{0}(k)$ then
10                   Set $M_{ij}=1$
11            else
12                   Set $M_{ij}=0$
13
14      Apply local mask: $\mathcal{P}(g_{k},k)=M\odot g_{k}$  // Following Eq. (5)
15       Update $\delta^{u}\leftarrow\prod_{\|\delta\|\leq\rho_{u}}\left(\delta^{u}-\alpha_{u}\cdot\mathrm{sign}(\mathcal{P}(g_{k},k))\right)$
16       Update source model parameter $\theta$ based on minibatch $(t(x+\delta^{u}),y)$
Output: SALM noise generator $f^{\prime}_{\theta}$, SALM noise $\delta^{u}$
Algorithm 1 Training SALM generator and generating noise.
3.3 Sparsity-Aware Local Masking
@@
To address the bi-level optimization problem in Eq. (2), existing studies have proposed methods such as iterative generator training [16], target poisoning with a pre-train model [15]. Specifically, in this work, we adopt the iterative generator training framework, with the training termination condition being solely the training steps $M$. When the training step $M$ is sufficient, the noise will be continued optimizing to achieve better performance since there’s no accuracy stop condition like [25]. In each step of noise update, we employ the PGD [44] to solve the constrained minimization problem as follows:
@@
where $t$ is the current step in the training process, $\nabla_{\boldsymbol{x}}\mathcal{L}\left(f_{\theta}^{\prime}\left(x_{t}^{\prime}\right)\right)$ is the gradient of loss, $\Pi$ is a projection function that clips the noise back to the refined area around the original example $x$ when it goes beyond, and $\alpha_{u}$ is the step size.
@@
With projection $\mathcal{P}$ to refine the local region, the perturbation $\delta$ is crafted with a PGD [44] process. Given total PGD steps $K_{a}$, for each iteration $t\in[1,K_{a}]$ the noise is iteratively updated via:
@@
By purposefully reducing the range of perturbation, this approach enables the noise to cover specific feature regions rather than dispersing throughout the entire color space. This ensures that the noise does not expend effort in sparse regions unnecessarily, leading to better performance and efficiency. The overall framework and procedure are depicted in Fig. 1 and Alg. 1.
4 Experiments
4.1 Setup
Datasets. Our research involved extensive experiments on MedMNIST [63], which is a comprehensive collection of standardized biomedical images spanning 12 distinct datasets. MedMNIST includes the major modalities in medical imaging. This dataset showcases diverse data scales and a range of tasks. Furthermore, to more effectively assess our method’s performance on large, real-world datasets, we also do experiments on the 224x224 datasets from MedMNIST-v2 [64].
Model and Implementation Details. We selected the following well-known models to evaluate our method against benchmark approaches: VGG-11 [57], ResNet-18/50 [22] and DenseNet-121 [24]. For the SALM generator, we select RN-18 as the source model to generate noise. And we set the perturbation noise $\rho_{u}=8/255$ as default. We choose an SGD optimizer for both noise generation and training. Both of their weight decay are set to 5e-4, and momentum is set to 0.9. The initial learning rate is set to 0.1 with a decay rate of 0.1.
Baselines. Since there’s no previous method specifically designed for medical data, we compare our Sparsity-Aware Local Masking (SALM) with the existing SoTA methods in the general image domain. The baseline methods are as following: Target Adversarial Poisoning (TAP) [15], Error-Minimizing noise (EM) [25], Synthesized Perturbation (SP) [67]. None of these three methods impose restrictions on the range of pixels to be perturbed. TAP and EM both target the training loss and craft noise using the gradient information to trigger maximum error or trick models to overfit. SC, on the other hand, imposes hand-crafted linear-separable noise into the data, thereby leading the model to learn only simple noise-label correlation.
4.2 Effectiveness Analysis
Figure 2: Left: The learning curves of ResNet-18 trained on different protected data. Right: The effect of different $K$ values on clean test accuracy(%) for the four datasets.
@@ 4.2 Effectiveness Analysis
PathMNIST 90.7 8.4(normal-$\downarrow$82.3) 13.5($\downarrow$77.2) 12.2($\downarrow$78.5) 11.8($\downarrow$78.9)
ChestMNIST 94.7 38.8($\downarrow$55.9) 78.7($\downarrow$16.0) 63.5($\downarrow$31.2) 27.6(normal-$\downarrow$67.1)
DermaMNIST 73.5 38.7($\downarrow$34.8) 17.1($\downarrow$56.4) 11.5($\downarrow$62.0) 8.7(normal-$\downarrow$64.8)
OCTMNIST 74.3 21.4($\downarrow$52.9) 25.0($\downarrow$49.3) 22.7($\downarrow$51.5) 21.2(normal-$\downarrow$53.1)
PneumoniaMNIST 85.4 66.9($\downarrow$18.5) 65.8($\downarrow$19.6) 70.1($\downarrow$15.3) 63.9(normal-$\downarrow$21.5)
RetinaMNIST 52.4 39.5($\downarrow$12.9) 13.0($\downarrow$39.4) 8.3(normal-$\downarrow$44.1) 12.2($\downarrow$40.2)
BreastMNIST 86.3 44.9(normal-$\downarrow$41.4) 46.2($\downarrow$40.1) 50.0($\downarrow$36.3) 53.2($\downarrow$33.1)
BloodMNIST 95.8 19.6($\downarrow$76.2) 30.5($\downarrow$65.3) 30.5($\downarrow$65.3) 7.1(normal-$\downarrow$88.7)
TissueMNIST 67.6 19.7($\downarrow$47.9) 17.4($\downarrow$50.2) 7.3(normal-$\downarrow$60.3) 21.4($\downarrow$46.2)
OrganAMNIST 93.5 81.0($\downarrow$12.5) 78.1($\downarrow$15.4) 86.1($\downarrow$7.4) 69.0(normal-$\downarrow$24.5)
OrganCMNIST 90.0 70.3($\downarrow$19.7) 72.2($\downarrow$17.8) 74.8($\downarrow$15.2) 64.6(normal-$\downarrow$25.4)
OrganSMNIST 78.2 51.7($\downarrow$26.5) 50.1($\downarrow$28.1) 53.0($\downarrow$25.2) 46.8(normal-$\downarrow$31.4)
[…]
We initially selected PathMNIST for comparison between our method and the conceptually similar error-minimizing and error-maximizing noise. The results in Fig. 2 left show that from the onset of training, both our method and the error-minimizing noise provide effective protection, with our method demonstrating superior performance. Conversely, the efficacy of error-maximizing noise gradually diminished throughout the training. To further evaluate the performance of our method, we directly transferred several conventional image protection methods to medical images and compared them with our approach. The results in Table 1 confirm that our method is not only broadly applicable to different modalities of medical datasets but also outperforms the previous SoTA methods for generic image protection.
4.3 Stability Analysis
@@
[…]
$\mathcal{D}_{c}$
$\mathcal{D}_{u}$
$\mathcal{D}_{c}$
$\mathcal{D}_{u}$
$\mathcal{D}_{c}$
$\mathcal{D}_{u}$
$\mathcal{D}_{c}$
$\mathcal{D}_{u}$
$K=10$ VGG-11
91.7
19.2
73.9
10.9
78.2
21.4
91.7
27.3
RN-18
90.7
11.8
73.5
8.7
74.3
21.2
92.7
25.4
RN-50
94.6
21.9
74.6
9.8
74.3
19.0
94.7
25.4
DN-121
96.4
14.8
75.3
14.5
75.6
13.3
95.1
21.2
$K=30$ VGG-11
90.8
18.3
74.4
11.8
75.5
15.1
92.1
31.4
RN-18
92.9
14.1
73.5
4.0
74.4
12.2
94.6
24.0
RN-50
94.2
17.4
74.4
9.3
77.0
14.3
94.8
30.7
DN-121
94.2
19.7
77.2
11.8
74.9
30.0
94.5
29.9
Different Architectures Selection. Before releasing datasets, we cannot foresee the specific training model unauthorized users might adopt. Hence, ensuring that the protected data remains effective across various models is crucial. While protectors can only select a single source model for noise generation, but unauthorized users are free to choose any model, seemingly placing the former at a disadvantage. However, in reality, the choice of model by users does not impact the efficacy of our method. The results in Table 2 demonstrate that the SALM with the source model ResNet-18 is effective across various models. Furthermore, the protection performance is not necessarily best when unauthorized trained on ResNet-18. This indicates that our method is not limited by the source model and the target model, enabling more effective application in real-world scenarios and facilitating the open sourcing of high-quality datasets. What’s more, the protective effect when $K$ is 30 is not significantly better than when it is 10. This prompts us to delve deeper into the impact of the choice of $K$-value.
Different $K$ Selection. In this section, we delve deeper into our approach to unveil the nuances of our method’s dependency on the choice of $K$ value. $K$ values were selected within a range from 0 to 90 in increments of ten. And we also tried $K=5$ to test the limitation of SALM. When $K$ is set to 0, it implies that the model is trained on clean data. The results in Fig. 2 right show when $K$ is set at 5 or higher, the noise proves to be generally effective. Consequently, in subsequent sections, unless specifically emphasized, we opt for $K=10$ for our experiments. This finding reinforces our assumption that perturbing only a few crucial pixels is enough. What’s more, it highlights the high efficiency of our approach to medical data, which can enable the accelerated release of datasets.
5 Sensibility Analysis

### v2

@@ Abstract
The rapid expansion of AI in healthcare has led to a surge in medical data generation and storage, boosting medical AI development. However, fears of unauthorized use, like training commercial AI models, hinder researchers from sharing their valuable datasets. To encourage data sharing, one promising solution is to introduce imperceptible noise into the data. This method aims to safeguard the data against unauthorized training by inducing degradation in the generalization ability of the trained model. However, they are not effective and efficient when applied to medical data, mainly due to the ignorance of the sparse nature of medical images. To address this problem, we propose the Sparsity-Aware Local Masking (SALM) method, a novel approach that selectively perturbs significant pixel regions rather than the entire image as previously. This simple yet effective approach, by focusing on local areas, significantly narrows down the search space for disturbances and fully leverages the characteristics of sparsity. Our extensive experiments across various datasets and model architectures demonstrate that SALM effectively prevents unauthorized training of different models and outperforms previous SoTA data protection methods.
Machine Learning, ICML
@@ 1 Introduction
The expansion of artificial intelligence (AI) in healthcare has significantly increased the production and storage of sensitive medical data (Rajkomar et al., 2018; Rasmy et al., 2021). This data plays a crucial role in advancing research in related fields (Zhang et al., 2022; Kelly et al., 2019; Rasmy et al., 2021). The more high-quality, open-source datasets that are available, the more contributions can be made by talented researchers to the development of the field (Johnson et al., 2019; Irvin et al., 2019; Pedrosa et al., 2019; Bilic et al., 2023; Rasmy et al., 2021). However, many dataset creators are reluctant to open-source their work due to concerns over unauthorized use, such as training commercial models (Edwards, 2022; Hill & Krolik, 2019).
@@
Conventional image classification datasets are relatively easy to obtain and label by laypersons, as shown by ImageNet’s use of online crowdsourcing (Deng et al., 2009; Zhang et al., 2021). But in the medical field, annotating is a much more complex process requiring specialized knowledge, usually undertaken by experts like radiologists (Kermany et al., 2018; Simpson et al., 2019; Kather et al., 2019). In addition, it is also necessary to describe the severity of the disease and its relationship with the surrounding tissues. This means that constructing a medical dataset is a labor-intensive task. Unauthorized use could lead to infringement of the creator’s rights, leading to a reluctance to release further data publicly. Moreover, from a patient’s perspective, concerns about their information being exploited for commercial purposes might decrease their willingness to authorize their data for research (Koh et al., 2011; Trinidad et al., 2020). Therefore, the construction and release of a medical dataset are not only time-consuming and labor-intensive but also fraught with significant ethical and privacy challenges. The reduction in quality open source datasets, resulting from both of the scenarios mentioned, in turn, slows down the development of medical AI (Alberdi et al., 2016; Forghani et al., 2019; Gunraj et al., 2020).
@@
To defend unauthorized use and encourage sharing data, Huang et al. (2021) proposed the technique of contaminating data with imperceptible noise. Models trained on this noise-contaminated data exhibit poor performance for normal utilization. Specifically, these methods create “unlearnable” examples from clean data by adding imperceptible noise, demonstrating significant protective capabilities. The design of this error-minimizing noise is based on the intuitive idea that an example with a higher training loss could contain more information to learn. Consequently, this noise protects the data by minimizing the corresponding loss, effectively reducing the informativeness of the data.
@@
Nevertheless, applying this method directly to medical images may not be optimal, as it overlooks the unique properties inherent in medical images (Liu et al., 2023d). The most important property that is overlooked is sparsity (Ye & Liu, 2012; Chuang et al., 2007; Huang et al., 2009; Otazo et al., 2015; Davoudi et al., 2019; Fang et al., 2013). For instance, in cellular microscopy images, even after cropping, there remains a substantial amount of blank background. Similarly, techniques such as CT or tomographic scanning inherently produce sparse data (Chen et al., 2018; Davoudi et al., 2019; Fang et al., 2013). Previous methods often struggled to pinpoint specific feature regions in medical data, inadvertently emphasizing sparse areas when generating noise. This led to a major waste of computational resources and suboptimal noise performance, impacts on protection effectiveness, and the time consumed for protection.
@@
    To address these issues, we propose Sparsity-Aware Local Masking (SALM), specifically designed for the medical domain by limiting the perturbation to features improved protection effectiveness and robustness.
@@
    Experiments on multiple medical datasets across different modalities, scales, and tasks demonstrate that our SALM achieves SoTA performance and is consistently effective in various medical scenarios.
@@ 2 Related Works
In this paper, we seek to protect medical data from unauthorized exploitation via a data poisoning approach. Data Poisoning is a technique used to compromise the performance of machine learning models on clean data by deliberately altering training samples. This form of attack has proven effective against both Deep Neural Networks (DNNs) and traditional machine learning methods such as SVM (Biggio et al., 2012). Muñoz-González et al. (2017) has highlighted the susceptibility of DNNs to data poisoning, although these attacks typically result in only a modest reduction in DNN performance. However, Yang et al. (2017) found there is a clear distinction between poisoned samples and normal samples, which comes at the cost of reduced data usability. Backdoor Attacks represent a specialized form of data poisoning. Traditional backdoor attacks (Chen et al., 2017; Liu et al., 2020) involve the introduction of falsely labeled training samples embedded with covert triggers into the dataset. A relatively new approach within this realm is the creation of Unlearnable Examples (Huang et al., 2021). These are considered a more subtle form of backdoor attack, free of labels and triggers. Unlearnable Examples show promising results in protecting data from unauthorized exploitation in various domains and applications (Liu et al., 2023c, b, a; Sun et al., 2022; Zhang et al., 2023; Li et al., 2023; He et al., 2022; Zhao & Lao, 2022; Salman et al., 2023; Guo et al., 2023), such as natural language processing (Ji et al., 2022), graph learning (Liu et al., 2023a), and contrastive learning (Ren et al., 2022). However, the flexibility of protecting medical data has not yet been fully explored. Recently, Liu et al. (2023d) took a first step to evaluate the performance of conventional image protection methods on medical images. Nevertheless, the methods in (Liu et al., 2023d) are straightforward adaptations from the previous approach and are suboptimal due to a lack of consideration of the intrinsic characteristics of medical data. To better address this, in this study, we leverage the inherent features of medical data to design more effective protection.
@@ 3.1 Problem Statement
Suppose the clean training dataset consists of $n$ clean examples, that is, $\mathcal{D}_{c}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{n}$ with $\boldsymbol{x}\in\mathcal{X}\subset\mathbb{R}^{d}$ are the inputs and $y\in\mathcal{Y}=\{1,\cdots,N\}$ are the labels and $N$ is the total number of classes. We denote its unlearnable version by $\mathcal{D}_{u}=\left\{\left(\boldsymbol{x}_{i}^{\prime},y_{i}\right)\right\}_{i=1}^{n}$, where $\boldsymbol{x}^{\prime}=\boldsymbol{x}+\boldsymbol{\delta}$ is the unlearnable version of training example $\boldsymbol{x}\in\mathcal{D}_{c}$ ,and $\boldsymbol{\delta}\in\Delta\subset\mathbb{R}^{d}$ is the “invisible” noise that makes $\boldsymbol{x}$ unlearnable. The noise $\boldsymbol{\delta}$ is bounded by $\|\delta\|_{p}\leq\epsilon$ with $\|\cdot\|_{p}$($L_{p}$ norm), and $\epsilon$ is set to be small such that it does not affect the normal utility of the example.
@@
3.2 Sparsity-Aware Local Masking
In an ideal scenario, the generation of noise involves a class-matching process to protect each category within $\mathcal{D}_{c}$. For simplicity, we define the noise generation process on $\mathcal{D}_{c}$ here. Given a clean example $\boldsymbol{x}$, by generating the imperceptible noise $\boldsymbol{\Delta}$ for the training input $\boldsymbol{x}$ by solving the following bi-level optimization problem:
@@ 3.2 Sparsity-Aware Local Masking
To address the bi-level optimization problem in Eq. (2), existing studies have proposed methods such as iterative generator training (Fu et al., 2022), target poisoning with a pretrained model (Fowl et al., 2021). Specifically, in this work, we adopt the iterative generator training framework, with the training termination condition being solely the training steps $M$. When the training step $M$ is sufficient, the noise will be continued optimizing to achieve better performance since there’s no accuracy stop condition like (Huang et al., 2021). In each step of noise update, we employ the PGD (Madry et al., 2017) to solve the constrained minimization problem as follows:
@@
where $t$ is the current step in the training process, $\nabla_{\boldsymbol{x}}\mathcal{L}\left(f_{\theta}^{\prime}\left(x_{t}^{\prime}\right)\right)$ is the gradient of loss, $\Pi$ is a projection function that clips the noise back to the refined area around the original example $x$ when it goes beyond, $\alpha_{u}$ is the step size.
@@
With projection $\mathcal{P}$ to refine the local region, the perturbation $\delta$ is crafted with a PGD (Madry et al., 2017) process. Given total PGD steps $K_{a}$, for each iteration $t\in[1,K_{a}]$ the noise is iteratively updated via:
@@
By purposefully reducing the range of perturbation, this approach enables the noise to cover specific feature regions rather than dispersing throughout the entire color space. This ensures that the noise does not expend effort in sparse regions unnecessarily, leading to better performance and efficiency. The overall framework and procedure are depicted in Figure 1 and Appendix B.
Figure 2: The learning curves of ResNet-18 trained on different protected data.
Figure 3: The selected categories protect effectiveness under different models.
Figure 4: The effect of $K$ on clean test accuracy(%) for the four datasets.
4 Experiments and Results
We selected more than 14 datasets from Medmnist (Yang et al., 2021) and Medmnist-v2 (Yang et al., 2023) and conducted extensive experiments on various models following (Huang et al., 2021). We compare SALM with three baselines, including AdvT (Fowl et al., 2021), EM (Huang et al., 2021), SP (Yu et al., 2022). For more experimental details, please refer to Appendix C. More experiments, analysis, and the case study can be found in Appendix 4 and E.
@@ 4 Experiments and Results
PathMNIST 90.7 8.4($\downarrow$82.3) 13.5($\downarrow$77.2) 12.2($\downarrow$78.5) 11.8($\downarrow$78.9)
ChestMNIST 94.7 38.8($\downarrow$55.9) 78.7($\downarrow$16.0) 63.5($\downarrow$31.2) 27.6($\downarrow$67.1)
DermaMNIST 73.5 38.7($\downarrow$34.8) 17.1($\downarrow$56.4) 11.5($\downarrow$62.0) 8.7($\downarrow$64.8)
OCTMNIST 74.3 21.4($\downarrow$52.9) 25.0($\downarrow$49.3) 22.7($\downarrow$51.5) 21.2($\downarrow$53.1)
PneumoniaMNIST 85.4 66.9($\downarrow$18.5) 65.8($\downarrow$19.6) 70.1($\downarrow$15.3) 63.9($\downarrow$21.5)
RetinaMNIST 52.4 39.5($\downarrow$12.9) 13.0($\downarrow$39.4) 8.3($\downarrow$44.1) 12.2($\downarrow$40.2)
BreastMNIST 86.3 44.9($\downarrow$41.4) 46.2($\downarrow$40.1) 50.0($\downarrow$36.3) 53.2($\downarrow$33.1)
BloodMNIST 95.8 19.6($\downarrow$76.2) 30.5($\downarrow$65.3) 30.5($\downarrow$65.3) 7.1($\downarrow$88.7)
TissueMNIST 67.6 19.7($\downarrow$47.9) 17.4($\downarrow$50.2) 7.3($\downarrow$60.3) 21.4($\downarrow$46.2)
OrganAMNIST 93.5 81.0($\downarrow$12.5) 78.1($\downarrow$15.4) 86.1($\downarrow$7.4) 69.0($\downarrow$24.5)
OrganCMNIST 90.0 70.3($\downarrow$19.7) 72.2($\downarrow$17.8) 74.8($\downarrow$15.2) 64.6($\downarrow$25.4)
OrganSMNIST 78.2 51.7($\downarrow$26.5) 50.1($\downarrow$28.1) 53.0($\downarrow$25.2) 46.8($\downarrow$31.4)
PathMNIST224 90.9 70.8($\downarrow$21.9) 32.5($\downarrow$60.2) 29.1($\downarrow$63.6) 27.5($\downarrow$67.3)
ChestMNIST224 92.7 31.6($\downarrow$61.1) 64.4($\downarrow$28.3) 50.1($\downarrow$42.6) 25.4($\downarrow$67.3)
Effectiveness Analysis. We initially selected PathMNIST for comparison between our method and the conceptually similar error-minimizing and error-maximizing noise. The results in Figure 4 show that from the onset of training, both our method and the error-minimizing noise provide effective protection, with our method demonstrating superior performance. Conversely, the efficacy of error-maximizing noise gradually diminished throughout the training. To further evaluate the performance of our method, we directly transferred several conventional image protection methods to medical and compared them with our approach. The results in Table 1 confirm that our method is not only broadly applicable to different modalities of medical datasets but also outperforms the previous SoTA methods.
@@
$K$ Model PathMNIST DermaMNIST OctMNIST ChestMNIST224
$\mathcal{D}_{c}$ $\mathcal{D}_{u}$ $\mathcal{D}_{c}$ $\mathcal{D}_{u}$ $\mathcal{D}_{c}$ $\mathcal{D}_{u}$ $\mathcal{D}_{c}$ $\mathcal{D}_{u}$
$K=10$ VGG-11 91.7 19.2 73.9 10.9 78.2 21.4 91.7 27.3
RN-18 90.7 11.8 73.5 8.7 74.3 21.2 92.7 25.4
RN-50 94.6 21.9 74.6 9.8 74.3 19.0 94.7 25.4
DN-121 96.4 14.8 75.3 14.5 75.6 13.3 95.1 21.2
$K=30$ VGG-11 90.8 18.3 74.4 11.8 75.5 15.1 92.1 31.4
RN-18 92.9 14.1 73.5 4.0 74.4 12.2 94.6 24.0
RN-50 94.2 17.4 74.4 9.3 77.0 14.3 94.8 30.7
DN-121 94.2 19.7 77.2 11.8 74.9 30.0 94.5 29.9
Different Architectures Selection. Before releasing datasets, we cannot foresee the specific training model unauthorized users might adopt. Hence, ensuring that the protected data remains effective across various models is crucial. While protectors can only select a single source model for noise generation, unauthorized users are free to choose any model, seemingly placing the protectors at a disadvantage place. However, in reality, the choice of model by unauthorized users does not impact the efficacy of our method. The results in Table 2 demonstrate that the SALM with the source model ResNet-18 (He et al., 2016) is effective across various models. The results in Figure 4 show the randomly extracted subset still has excellent performance under different models. Furthermore, the protection performance is not necessarily best when unauthorized trained on ResNet-18. This indicates that our method is not limited by the source model and the target model, enabling more effective application in real-world scenarios and facilitating the open-sourcing of high-quality datasets. What’s more, the protective effect when $K$ is 30 is not significantly better than when it is 10. This prompts us to delve deeper into the impact of the choice of $K$-value.
Different $K$ Selection. In this section, we delve deeper into our approach to unveil the nuances of our method’s dependency on the choice of $K$ value. $K$ values were selected within a range from 0 to 90 in increments of ten. We also tried $K=5$ to test the limitation of SALM. When $K$ is set to 0, it implies that the model is trained on clean data. The results in Figure 4 show when $K$ is set at 5 or higher, the noise proves to be generally effective. Consequently, in subsequent sections, unless specifically emphasized, we opt for $K=10$ for our experiments. This underscores the effectiveness of limiting the perturbation search space and also demonstrates that the existing methods are not optimal in the medical domain. What’s more, it highlights the high efficiency of our approach to medical data, which can enable the accelerated release of datasets.